Let m > 0 be a constant. Suppose a tank starts with 10kg of salt water with 1% salt by mass. Suppose further that every hour,(b) (5 points) How long does it take for the water to reach 5% concentration if m= 3.(5 points) Give a (simplified) equation describing a value for m such that the solution to reaches exactly 5% concentration i
Answer #1

Please give LIKE

sont from the given data mo- Constant m mars lokg of salt water ly. salt by mas - Concentication of salt 10% cidded. akg of wa .. The differential equation for the given data is given by dult = m - X(t) at 10 10+(m-2)= dxCt) + 2(E) Em 10 10+ m2) E le+C (0+t) )=S. C10H8) dt +c (to++) 260) = Rot +)+c xlt) = 3+ (20+t) + E 20 (10+4) Ciort) Y at t=0, alo) = 1 gives c=1. is the(10++)?= 6ot + 3t+20 100+206 +ť = 3t + 6ot +20 34_t? + Get-aot +20-400 = 0 at +40t –80 =0 ť +20t -40 =0 Gires ta -218 & 1.83